For what values of k does the equation 2x squared + (1-k)x+8=0 have a double root? I have tried numerous ideas, but cannot figure this out.
Answer by scott8148(6628)

a double root means a perfect square

dividing by 2 gives x^2+((1-k)/2)x+4=0

taking the square roots of 4 leads us to (x+2)^2=x^2+4x+4 and (x-2)^2=x^2-4x+4

(1-k)/2=4 ___ 1-k=8

(1-k)/2=-4 ___ 1-k=-8